OpenStudy (anonymous):

Explain why csc 180° is undefined

OpenStudy (anx):

csc is the same as 1/sin so on the unit circle sin 180 = 0 so csc = 1/sin = 1/sin(180) - 1/ 0 which is undefined.

OpenStudy (anonymous):

csc 180=1/sin 180. sin 180= 0, anything when divided by 0 is undefined. hence scs 180 is undefined.@mackenziecarmody37

OpenStudy (solomonzelman):

cosecant and secant are undefined for this reason. It is just like having a variable with negative exponent equal to zero.